OpenStudy (anonymous):

how do I solve (k+1)(k-5)=0 quadratic by factoring

OpenStudy (anonymous):

@EvelyngLopez are you still looking for help?

OpenStudy (clalgee):

Welcome to OpenStudy @EvelyingLopez ! Since the equation is already factored. All you need to do now is use the zero product property. (k + 1) (k - 5)= 0 k + 1= 0 or k - 5= 0 k= -1 or k= 5

OpenStudy (clalgee):

@EvelyngLopez